### v1

# Sequential Monte Carlo Squared for Online Inference in Stochastic Epidemic Models

Abstract
Effective epidemic modeling and surveillance require computationally efficient methods that can continuously update estimates as new data becomes available. This paper explores the application of an online variant of Sequential Monte Carlo Squared (O-SMC2) to the stochastic Susceptible-Exposed-Infectious-Removed (SEIR) model for real-time epidemic tracking. The particularity of O-SMC2 lies in its ability to update the parameters using a particle Metropolis-Hastings kernel, ensuring that the target distribution remains invariant while only utilizing a fixed window of recent observations. This feature enables timely parameter updates and significantly enhances computational efficiency compared to the standard SMC2, which processes the entire dataset. First, we demonstrate the efficiency of O-SMC2 on simulated data, where both the parameters and the observation process are known. We then apply the method to a real-world COVID-19 dataset from Ireland, successfully tracking the epidemic trajectory and estimating the time-dependent reproduction number of the disease. Our results show that O-SMC2 provides highly accurate online estimates of both static and dynamic epidemiological parameters while substantially reducing computational costs. These findings highlight the potential of O-SMC2 for real-time epidemic monitoring and supporting adaptive public health interventions.
Keywords: Disease modeling, Stochastic model, Sequential Monte Carlo, Online inference
1 Introduction
Disease modeling plays a crucial role in public health by providing insights into the mechanisms behind the spread of infectious diseases. Such understanding is vital for designing effective intervention strategies and supporting evidence-based policy-making. As the global landscape of infectious diseases evolves and the risk of novel outbreaks increases (Marani et al., 2021), there is a pressing need for robust modeling frameworks that can adapt to changing data, update parameter values as new information becomes available, and deliver accurate assessments and predictions of epidemic trajectories (Birrell et al., 2017).
Bayesian approaches have emerged as powerful tools for tackling these challenges, particularly in sampling from the posterior distribution of model parameters. Markov chain Monte Carlo (MCMC) methods have been widely utilized in epidemic analysis (O’Neill and Roberts, 1999; Jewell et al., 2009; Wang and Walker, 2022), enabling exact inference but requiring an explicit expression for the complete-data likelihood. Furthermore, these methods can be computationally intensive and inefficient for real-time applications, as they necessitate reanalyzing the entire dataset with each new data point (Birrell et al., 2017). An alternative approach is Approximate Bayesian Computation (ABC), which has gained popularity for analyzing epidemic processes due to its ability to bypass the need for explicit likelihood formulations (Kypraios et al., 2017). While ABC offers flexibility, it relies on summary statistics, which can introduce bias and limit its precision (Kypraios et al., 2017). Furthermore, ABC does not inherently address the challenge of sequentially updating estimates as new data are collected.
Sequential Monte Carlo (SMC) methods, also known as particle filtering, offer an attractive solution to update the posterior distribution of a stochastic process as new data become available (Doucet et al., 2001). These methods have seen increasing application in epidemiology, with early studies demonstrating their effectiveness in calibrating nonlinear compartmental infectious disease models (Yang et al., 2014; Camacho et al., 2015). For instance, Welding and Neal (2019) employed a sequential Bayesian framework to the 2001 UK Foot-and-Mouth epidemic, utilizing MCMC algorithms to rejuvenate particles and facilitate real-time outbreak analysis through data augmentation. Additionally, Birrell et al. (2020) proposed an age-structured SEIR model for influenza outbreaks, showcasing the advantages of SMC over MCMC for online prediction, particularly in integrating multiple data sources for parameter inference.
However, it is important to recognize that traditional SMC algorithms are not well suited for estimating static parameters (Kantas et al., 2015). Although extensions like particle MCMC (PMCMC) (Andrieu et al., 2010) have been developed to achieve exact sampling from the posterior by creating unbiased estimators for the likelihood through simulation and subsequently applying traditional MCMC methods, they remain computationally demanding. Each MCMC iteration typically requires a full SMC run, making these methods less suitable for real-time applications (Dureau et al., 2013; Funk et al., 2018; Endo et al., 2019; Cazelles et al., 2021). In their recent study, Storvik et al. (2023) propose an SMC framework for inferring a time-varying reproduction number of COVID-19 in Norway. It is essential to highlight that this SMC algorithm relies on a set of sufficient statistics to sequentially estimate hyperparameters that control the evolution of the reproduction number. However, the estimation of the posterior parameter relies on the assumption that sufficient statistics-based filters are analytically tractable, which will only hold for particular dynamical system setups.
A promising solution to these challenges is the SMC2 method (Chopin et al., 2013; Jacob, 2015), which enhances traditional particle filtering by integrating a nested filtering mechanism. This approach simultaneously estimates model states and parameters, improving both the accuracy and adaptability of the estimation process. A key feature of the SMC2 algorithm is the “rejuvenation” step, which uses a PMCMC kernel to mitigate particle impoverishment. However, this step traditionally requires computing the likelihood using all past data points, which may limit its online applicability.
In this paper, we implement an online variant of the SMC2 method (O-SMC2), which reduces computational costs per iteration by targeting an approximate distribution and evaluating the PMCMC over a fixed window of recent observations (Vieira, 2018). This adaptation preserves the sequential and online attributes of the method, emphasizing its practical application in real-time state and parameter estimation for epidemic models, specifically the stochastic SEIR model.
The remainder of this paper is structured as follows. Section 2 describes the state-space framework for the SEIR compartmental model and provides a brief overview of the SMC2/O-SMC2 methodology. In Section 3, we apply the O-SMC2 algorithm to both simulated data and real-world COVID-19 data from Ireland, demonstrating its potential for real-time epidemiological surveillance and for inferring the time-dependent reproduction number. Finally, Section 4 concludes the paper by summarizing the main findings and discussing the potential challenges and limitations associated with our model.
2 Compartmental model and sequential inference
2.1 Stochastic SEIR model
Compartmental models for infectious diseases, which had their conceptual origins in the early to mid $20^{th}$ century, attempt to describe a disease’s progression and spread through individuals moving through disease stages over time. The stochastic SEIR model serves as the foundation for our investigation. Consider the time interval $(t,t+\delta t]$, where $\delta t$ represents the time between observations. If we assume that the time spent by an individual in a compartment is exponentially distributed with a rate $\lambda$, then the number of individuals leaving a compartment during the time step $\delta t$ can be seen as a Binomial variable with the probability $p=1-e^{-\lambda\delta t}$. This parameterization ensures that the probability will remain between 00 and $1$. Hence the discrete-time SEIR stochastic model is given by:
$S_{t+\delta t}$ $=S_{t}-Y_{SE}(t),$ $Y_{SE}(t)$ $\sim\mathrm{Bin}\left(S_{t},1-e^{-\beta_{t}\frac{I_{t}}{N}\delta t}\right)$
$E_{t+\delta t}$ $=E_{t}+Y_{SE}(t)-Y_{EI}(t),$ $Y_{EI}(t)$ $\sim\mathrm{Bin}\left(E_{t},1-e^{-\sigma\delta t}\right)$ (1)
$I_{t+\delta t}$ $=I_{t}+Y_{EI}(t)-Y_{IR}(t),$ $Y_{IR}(t)$ $\sim\mathrm{Bin}\left(I_{t},1-e^{-\gamma\delta t}\right)$
$R_{t+\delta t}$ $=R_{t}+Y_{IR}(t).$
Here, $S_{t}$, $E_{t}$, $I_{t}$, and $R_{t}$ represent the compartments of susceptible, exposed (infected but not yet infectious), infected, and removed individuals, respectively, at time $t$. Susceptible individuals, upon contact with infected individuals, transition to the exposed class at a rate $\beta_{t}$ (which can be time-dependent or constant). Exposed individuals become infectious at a rate $\sigma$, and infected individuals recover from the disease at a rate $\gamma$. The population is assumed to have a constant size $N=S_{t}+E_{t}+I_{t}+R_{t}$ at each time point.
Figure 1: Graphical representation of a SEIR model. Each arrow expresses the rate of transition between the compartments. $\beta_{t}$ is the disease transmission rate, $1/\sigma$ is the mean incubation period, and $1/\gamma$ is the mean infectious period.
2.2 State-space framework
In the context of the SEIR model, the system can be formulated as a state-space model (SSM), which consists of two stochastic processes: the first process governs the evolution of the disease within the population, while the second process represents the observation process. The state-space model is defined for time steps $t=1,\ldots,T$ as follows:
[…]
Here, $x_{t}$ denotes the latent state variables (state variable in the classical SEIR model), which evolve according to a Markov process with transition distribution $f(x_{t}|x_{t-1},\theta)$ given by the model (2.1). The observation $y_{t}$ is assumed to be probabilistically related to the hidden state $x_{t}$ via the distribution $g(y_{t}|x_{t},\theta)$. The parameter vector $\theta$ encompasses all static epidemiological parameters of the model such as $\beta$, $\sigma$ and $\gamma$, governing both the state transitions and the observation process. Throughout the study, $\theta$ will be used interchangeably to refer to these parameters.
Figure 2: Graphical representation of a state-space model. Black arrows show variable dependencies, and dashed gray lines represent the dependence on $\theta$.
State-space models are particularly useful for modeling complex dynamical systems such as disease transmission where the true states are not directly observable, and observations are subject to noise or other forms of uncertainty. By explicitly modeling the hidden states and their relationship to observed data, we can perform inference on both the state variables and the model parameters.
2.3 State and Parameter inference
In this section, we briefly describe the procedure for sequential parameter estimation using the SMC2 scheme. Let $x_{0:t}$  [^1] For any sequence $\{z_{t}\}_{t\geq 0}$, we denote $z_{i:j}=\{z_{i},z_{i+1},\ldots,z_{j}\}$ and $z^{1:N}_{j}$ (resp. $z^{1:N}_{i:j}$) is a set of $N$ random variables $z^{l}_{j}$ (resp. $z^{l}_{i:j}$), for $l=1,...,N$. represent the sequence of latent states (in our case the SEIR compartment counts) up to time $t$, and let $y_{1:t}$ represent the observations (e.g., daily reported cases). The goal is to construct a sequence of posterior distributions $\{p(\theta|y_{1:t}),t=1,\ldots,T\}$ that reflect the evolution of parameter estimates as new observations become available. Bayes’ rule provides the basis for updating the posterior distribution as follows:
$p(\theta|y_{1:t})\propto p(\theta|y_{1:t-1})p(y_{t}|y_{1:t-1},\theta),$ (5)
which suggests a sequential importance sampling approach. This approach involves reweighting a set of parameter particles based on the incremental likelihood $p(y_{t}|y_{1:t-1},\theta)$. The iterated batch importance sampling (IBIS) method (Chopin, 2002) employs this technique with an MCMC step to rejuvenate parameter samples. However, for state-space models, the incremental likelihood is generally intractable. To address this issue, we can replace $p(y_{t}|y_{1:t-1},\theta)$ with an unbiased estimate derived from running a particle filter, targeting the filtering distribution $p(x_{t}|y_{1:t},\theta)$ for each of the $\theta$-particles. This approach is known as SMC2 due to the use of nested filters. For a comprehensive discussion on this approach, including formal justifications and recent developments, refer to (Chopin et al., 2013; Jacob, 2015).
2.3.1 Particle filtering
For a given value of the model parameters $\theta$, we aim to infer the latent variables $x_{t}$ at each time point, given the observations $y_{1:t}$ by means of the filtering distribution $p(x_{t}|y_{1:t},\theta)$. Particle filtering is a technique that approximates this filtering distribution by sequentially generating a set of particles with associated weights $\{x_{t}^{i},w_{t}^{i}\}_{i=1}^{N_{x}}$, via importance sampling and resampling. Hence, for a known parameter value $\theta$, the posterior distribution of the latent states is represented as a Dirac delta mixture:
$p(x_{t}|y_{1:t},\theta)\approx\dfrac{1}{\sum_{j=1}^{N_{x}}w_{t}^{j}}\sum_{i=1}^{N_{x}}w_{t}^{i}\delta_{x_{t}^{i}}(x_{t}).$ (6)
Here, $N_{x}$ denotes the number of $x$-particles, and $w^{i}_{t}:=w_{t}^{i}(x_{t-1}^{i},x_{t}^{i})$ represents the importance weight for each particle $x_{t}^{i}$. The weight update rule at time $t$ takes into account both the likelihood of the new observation $y_{t}$ given the current particle state $x_{t}^{i}$, and the transition probability between successive states:
$w_{t}^{i}=w_{t-1}^{i}\dfrac{g(y_{t}|x_{t}^{i},\theta)f(x_{t}^{i}|x_{t-1}^{i},\theta)}{q(x_{t}^{i}|x_{t-1}^{i},y_{t},\theta)},\quad i=1,\ldots,N_{x},$ (7)
One notable particle filtering method is the Bootstrap particle filter (BPF) (Gordon et al., 1993), which simplifies weight calculations by setting the importance distribution equal to the state transition distribution, i.e, $q(x_{t}|x_{t-1},y_{t},\theta)=f(x_{t}|x_{t-1},\theta)$. Other advanced filtering algorithms have been developed to improve efficiency (Pitt and Shephard, 1999), but these are not discussed in this study. The general procedure for the Bootstrap Particle Filter is shown in Algorithm 1.
Algorithm 1 Bootstrap Particle Filter (BPF) for a given parameter $\theta$
Operations involving index $i$ must be performed for $i=1,...,N_{x}$.
The indices $a_{t}^{1:N_{x}}$ define the ancestral particles at time $t$ after the resampling.
Inputs: Observation: $y_{1:T}$, Number of particles: $N_{x}$, Inital state distribution: $f(x_{0}|\theta)$, Parameter vector: $\theta$.
Output: Particles set :$\left\{x^{i}_{0:t},w^{i}_{0:t}\right\}_{i=1}^{N_{x}}$
1:Sample initial particles : $x^{i}_{0}\sim f(x_{0}|\theta)$
2:Compute weights: $w^{i}_{0}=1$,  $W^{i}_{0}=w^{i}_{0}/\sum_{j=1}^{N_{x}}w^{j}_{0}$
3:for $t=1$ to $T$ do
4:      Sample new indices: $a_{t}^{1:N_{x}}\sim\text{Resample}(W_{t-1}^{1:N_{x}})$
5:     Propagate states $x_{t}^{i}\sim f(\cdot|x_{t-1}^{a_{t}^{i}},\theta)$
6:     Compute new weights $w_{t}^{i}=g(y_{t}|x_{t}^{i},\theta)$ ,  $W^{i}_{t}=w^{i}_{t}/\sum_{j=1}^{N_{x}}w^{j}_{t}$
7:end for
The resampling step in Algorithm 1, referred to in step 4, aims to mitigate particle degeneracy [^2] Particle degeneracy refers to the situation where, after many iterations, most particle weights approach zero while only a few retain significant weight. by discarding particles with lower weights and duplicating those with higher weights. After resampling, we reset the weights to $w^{i}_{t-1}=1$. Various resampling schemes exist as described in (Doucet et al., 2009). Unless stated otherwise, we use stratified resampling in all our implementations due to its efficiency and lower computational cost (Sheinson et al., 2014; Vieira, 2018). In Algorithm 1, if the resampling step is omitted, then the weights are replaced by $w^{i}_{t}=w^{i}_{t-1}g(y_{t}|x_{t}^{i},\theta)$. Particle filtering can also be used for Bayesian parameter inference by providing an unbiased estimate of the incremental likelihood:
$p(y_{t}|y_{1:t-1},\theta)\approx\hat{p}_{N_{x}}(y_{t}|y_{1:t-1},\theta)=\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{t}^{i}$ (8)
Hence, by the law of total probability, we can deduce an SMC estimate of the marginal likelihood for a given parameter $\theta$:
$p(y_{1:t}|\theta)\approx\hat{p}_{N_{x}}(y_{1:t}|\theta)=\prod_{s=1}^{t}\hat{p}_{N_{x}}(y_{s}|y_{1:s-1},\theta)=\prod_{s=1}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}\right).$ (9)
2.3.2 SMC2 scheme
Given a weighted sample $\{\theta^{m}_{t},\omega^{m}_{t}\}_{m=1}^{N_{\theta}}$ from $p(\theta|y_{1:t})$, the SMC2 algorithm reweights each $\theta$-particles according to an unbiased estimate of the incremental likelihood $p(y_{t}|y_{1:t-1},\theta^{m}_{t})$, obtained from the particle filter. As time progresses, there is a risk that only a few $\theta$-particles will accumulate significant weight. To address this, a resampling step is performed to generate a new set of particles from the dominant ones, followed by a PMCMC move step, such as the particle Marginal Metropolis-Hastings (PMMH) kernel, which preserves the target posterior distribution (Andrieu et al., 2010). The combination of the resampling and move steps is referred to as the “rejuvenation” step. In this study, new parameters are proposed from a normal distribution $\theta^{*}\sim q(\theta^{*}|\theta^{m}_{t}):=\mathcal{N}(\hat{\mu}_{t},c\hat{\Sigma}_{t})$, where $c$ is a scaling factor. The mean and variance-covariance matrix are given by:
$\hat{\mu}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\theta^{m}_{t},\text{ and }\hat{\Sigma}_{t}=\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}(\theta^{m}_{t}-\hat{\mu}_{t})(\theta^{m}_{t}-\hat{\mu}_{t})^{T},$ (10)
This allows us to draw new parameter particles in the region where most of the probability mass is located (Chopin, 2002).
Algorithm 2 SMC2
Operations involving index $m$ must be performed for $m=1,...,N_{\theta}$.
The indices $a_{1:N_{\theta}}$ define the ancestral particles after the resampling.
Inputs: Observation: $y_{1:T}$, Prior: $p(\theta)$, Number of state particles: $N_{x}$, Number of parameter particles: $N_{\theta}$, Resample threshold : $\tau_{R}$, Move step: $M$
Output: Particle set: $\left\{\theta^{m}_{0:t},\omega^{m}_{0:t},x_{0:t,\theta^{m}}^{1:N_{x}},w_{0:t,\theta^{m}}^{1:N_{x}}\right\}_{m=1}^{N_{\theta}}$.
1:Sample a initial particles $\theta^{m}_{0}\sim p(\theta)$ and set $\omega^{m}_{0}=1/N_{\theta}$.
2:for $t=1$ to $T$ do
3:     Perform iteration $t$ of the BPF to obtain $\left\{x_{t,\theta^{m}_{t}}^{1:N_{x}},w_{t,\theta^{m}_{t}}^{1:N_{x}}\right\}$ and $\hat{p}_{N_{x}}(y_{t}|y_{1:t-1},\theta^{m}_{t})$.
4:     Update importance weights: $\omega^{m}_{t}=\omega^{m}_{t-1}\hat{p}_{N_{x}}(y_{t}|y_{1:t-1},\theta^{m}_{t}),\quad\omega^{m}_{t}=\omega^{m}_{t}/\sum_{j=1}^{N_{x}}\omega^{j}_{t}$
5:     if ESS $<\tau_{R}N_{\theta}$ then
6:         Sample new indices $a_{1:N_{\theta}}\sim\text{Resample}(\omega^{1:N_{\theta}}_{t})$
7:         Set $\left\{\theta^{m}_{t},\omega^{m}_{t}\right\}:=\left\{\theta_{t}^{a_{m}},1/N_{\theta}\right\}$, $\left\{x_{t,\theta^{m}_{t}}^{1:N_{x}},w_{t,\theta^{m}_{t}}^{1:N_{x}}\right\}:=\left\{x_{t,\theta_{t}^{a_{m}}}^{1:N_{x}},w_{t,\theta^{a_{m}}_{t}}^{1:N_{x}}\right\}$
8:          Propose $\theta^{*}\sim q(\theta^{*}|\theta^{m}_{t})$ and perform BPF with $N_{x}$ particles given $\theta^{*}$ to obtain $\hat{p}_{N_{x}}(y_{1:t}|\theta^{*})$
9:
10:         Compute : $\alpha=\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{1:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{1:t}|\theta^{m}_{t})p(\theta^{m}_{t})}\times\dfrac{q(\theta^{m}_{t}|\theta^{*})}{q(\theta^{*}|\theta^{m}_{t})}\right\}$
11:
12:          Set $\big{(}\theta_{t}^{m},x_{t,\theta^{m}_{t}}^{1:N_{x}},w_{t,\theta^{m}_{t}}^{1:N_{x}}\big{)}=\begin{cases}\big{(}\theta^{*},x_{t,\theta^{*}}^{1:N_{x}},w_{t,\theta^{*}}^{1:N_{x}}\big{)}\leavevmode\nobreak\ \text{ with probability }\leavevmode\nobreak\ \alpha\\ \big{(}\theta_{t}^{m},x_{t,\theta^{m}_{t}}^{1:N_{x}},w_{t,\theta^{m}_{t}}^{1:N_{x}}\big{)}\leavevmode\nobreak\ \text{ with probability }\leavevmode\nobreak\ 1-\alpha\end{cases}$
13:     end if
14:end for
The $\mathrm{SMC}^{2}$ scheme is outlined in Algorithm 2, and the model evidence can be estimated using the following expression:
$\hat{Z}_{t}=\prod_{s=1}^{t}\left(\dfrac{1}{\sum_{m=1}^{N_{\theta}}\omega^{m}_{s-1}}\sum_{m=1}^{N_{\theta}}\omega^{m}_{s-1}\hat{p}_{N_{x}}(y_{s}|y_{1:s-1},\theta^{m}_{s})\right),$ (11)
where $\hat{p}_{N_{x}}(y_{s}|y_{1:s-1},\theta^{m}_{s})$ is the likelihood computed using the particle filter, as described in equation (8). Each rejuvenation (step 8-12 of the Algorithm 2) performs $M$ successive PMMH moves on each particle $\theta^{m}_{t}$. The resampling step is similar to the one used in Algorithm 1 but is performed only when a degeneracy criterion is met. A common way is to estimate the degree of degeneracy by computing the effective sample size (ESS) :
$\text{ESS}=\left(\sum_{m=1}^{N_{\theta}}\omega^{m}_{t}\right)^{2}\bigg{/}\sum_{m=1}^{N_{\theta}}(\omega^{m}_{t})^{2}.$ (12)
The rejuvenation step is triggered if the ESS falls below a specified threshold, typically if ESS $<\tau_{R}N_{\theta}$, where $\tau_{R}\in(0,1)$ is a user-defined parameter. The choice of $\tau_{R}$ controls the frequency of rejuvenation steps. A lower threshold for ESS leads to more frequent rejuvenation, which increases the computational cost. In practice, an ESS threshold of 50% ($\tau_{R}=0.5$) is commonly used (Chopin et al., 2013; Jacob, 2015).
2.4 Online SMC2
While SMC2 allows for sequential estimation of parameters as new observations arrive, it is not a truly online method. The PMMH step, which has a computational complexity of $\mathcal{O}(tN_{\theta}N_{x})$, involves a calculation of the total likelihood. This complexity increases with the number of observations $t$, making it computationally expensive as more data becomes available. Although Chopin et al. (2013) demonstrates that, in theory, the frequency at which the rejuvenation step is performed will decrease over time, the algorithm may still be impractical for some real-time applications. A possible solution suggested by Vieira (2018) to implement SMC2 in an online manner (O-SMC2) is to restrict the likelihood calculation set to a fixed-size subset of past observations, keeping the computational cost bounded. By considering a fixed window of observations of size $t_{k}>0$ in the PMMH, we can ensure that the online part of SMC2 will not be altered, and the online approximation will target the posterior of the parameters given the observation window $p(\theta|y_{t-t_{k}:t})$. This approach changes the cost from $\mathcal{O}(tN_{\theta}N_{x})$ to $\mathcal{O}(t_{k}N_{\theta}N_{x})$.
Hence, in step 8 of Algorithm 2, once we have reached $t>t_{k}$, the Metropolis-Hastings acceptance probability will now take the form:
$\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{m}_{t})p(\theta^{m}_{t})}\times\dfrac{q(\theta^{m}_{t}|\theta^{*})}{q(\theta^{*}|\theta^{m}_{t})}\right\},$ (13)
where the estimate of the marginal likelihood within the time window is computed similarly to (9), but this time using the restricted particle set $x_{t-t_{k}:t}^{1:N_{x}}$ ( see A for further details) :
$\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right).$ (14)
This restriction is possible under the assumption that static parameters will remain constant indefinitely (Vieira, 2018). The O-SMC2 scheme appears to be particularly amenable to parallelization over parameter particles, as the rejuvenation step can be performed separately for each parameter value.
3 Experimental and Real-world data analysis
To illustrate the methodology described in the previous sections, we conduct a series of experiments using both simulated and real-world data. The algorithm is implemented in Python version 3.10.12 and was executed on a desktop computer with an Intel Core i7-1300H processor and a 3.40 GHz clock speed. The code is available at https://github.com/Dhorasso/smc2-seir-model.
3.1 Experimental setup
We begin by applying the filtering algorithm to two synthetic datasets generated using the stochastic SEIR model introduced in Section 3. The simulation period spans $T=100$ days, with a daily time step of $\delta t=1$ day. The algorithm utilizes $N_{\theta}=400$ parameter particles, $N_{x}=200$ state particles, with $M=5$ successive PMMH moves and a window size of $t_{k}=20$ days. In Experiment 1, all model parameters remain constant, while Experiment 2 introduces a time-varying transmission rate. We assume that the data on daily new cases are the realization of a Poisson distribution: $y_{t}|x_{t}\sim\mathrm{Poisson}(Y_{EI}(t))$, where $Y_{EI}(t)$ is the number of individual moving from $E$ to $I$ at time $t$.
3.1.1 Experiment 1
In this experiment, we assume constant parameters for the SEIR model: $\beta=0.6$, $\sigma=\frac{1}{3}$, and $\gamma=\frac{1}{5}$. The total population size is $N=6000$, with one initially infected individual. The latent state at time $t$ is represented by $x_{t}=(S_{t},E_{t},I_{t},R_{t})$, and the parameters $\theta=(\beta,\sigma,\gamma)$ are to be estimated. The initial particle state vector in the O-SMC2 algorithm is set to $(S_{0},E_{0},I_{0},R_{0})=(N-I_{0},0,I_{0},0)$, with $I_{0}$ drawn from a discrete uniform distribution, $I_{0}\sim\mathcal{U}({0,\dots,5})$. A uniform prior on the interval $[0,1]$ is assigned to each of the parameters $\beta$, $\sigma$, and $\gamma$.
Figure 3: State and parameter inference for Experiment 1. Top row: Filtering estimates of daily new infections, with observed values shown as orange dots. Middle row: Time series of filtering estimates for parameters $\beta$, $\sigma$, and $\gamma$. Bottom row: Posterior distributions of $\beta$, $\sigma$, and $\gamma$ at the final time step, with 95% credible intervals in dashed blue lines. In rows 1 and 2, the solid black lines show the median estimate; shaded blue regions indicate posterior uncertainty quantiles. True parameter values are represented by orange dashed lines.
Figure 3 displays the sequential inference of daily new cases of infection together with the static parameters. We can observe that the predicted new cases are in good agreement with the data. The results also indicate that all true parameter values fall within the 50% credible interval, closely aligning with the posterior median as we evolve over time. Initially, during the epidemic’s onset, the data lacks sufficient information within the time interval $[0,20]$, resulting in widespread parameter uncertainty. However, as we get closer to the peak of the epidemic, the inference algorithm captures the dynamics of the epidemic, notably reducing parameter uncertainty. As the dynamics begin to slow around time $t=60$, the posterior uncertainty stays about the same.
3.1.2 Experiment 2
In this experiment, we introduce variability in the transmission rate to simulate a more realistic scenario. Synthetic data are generated using a time-varying $\beta(t)$, modeled as:
$\beta(t)=0.3\exp\left(\sin\left(\frac{2\pi t}{55}\right)-\frac{t}{80}\right),$ (15)
while the parameters $\sigma=\frac{1}{2}$ and $\gamma=\frac{1}{7}$ remain constant. The total population size is $N=200,000$, with five initially infected individuals. This experiment assesses the ability of the O-SMC2 algorithm to capture state and static parameters under high transmission variability.
To accommodate the non-constant transmission rate, we model $\beta_{t}$ in the model (2.1) as a geometric Brownian motion:
$\beta_{t}=e^{\log(\beta_{t-1})+\varepsilon_{t}},\quad\varepsilon_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (16)
[…]
Figure 4: State and parameter inference for Experiment 2. Top row: Left - Filtering estimates of daily new infections with observed values in orange; Right - Filtering estimate of the time-dependent transmission rate $\beta_{t}$. Middle row: Filtering estimates of $\sigma$, $\gamma$, and $\nu_{\beta}$ over time. Bottom row: Posterior distributions of $\sigma$, $\gamma$, and $\nu_{\beta}$ the final time step, with 95% credible intervals in dashed blue lines. In rows 1 and 2, the solid black lines show the median estimate; shaded blue regions indicate posterior uncertainty quantiles. True parameter values are represented by orange dashed lines.
Figure 4 demonstrates the ability of the method to track multiple outbreak waves and capture the trend in a time-varying transmission rate, albeit with a minor lag. This slight delay is expected, as the analysis relies on daily case data (i.e., the number of newly infected individuals transitioning from $E$ to $I$) rather than exact counts of new infections (i.e., individuals moving from $S$ to $E$), which are often unknown during an epidemic. As with Experiment 1, the true parameter values lie within the 50% credible interval, indicating reliable parameter estimation even under conditions of variable transmission.
To evaluate the performance of the O-SMC2 algorithm, we tested a range of window sizes $t_{k}$, keeping the priors and number of particle configurations consistent across experiments. Tables B.1 and B.2 in C show that even with smaller window sizes, the algorithm produces reasonable parameter estimates that are close to the true values, with credible intervals comparable to those of standard SMC2 (although slightly larger for smaller window sizes). These findings align with previous research (Vieira (2018), Chapter 15). Figures B.1-B.2 display the posterior densities of the parameters at the final time step, demonstrating consistency with estimates obtained using SMC2.
3.2 Real case study: Analysis of COVID-19 in Ireland
Since the onset of the COVID-19 pandemic in December 2019 in Wuhan, China, the World Health Organization (WHO) has reported over 772 million confirmed cases globally and more than 69 million deaths (World Health Organisation (2024), WHO). The rapid transmission and widespread nature of the virus led the WHO to declare it a pandemic. In (Cazelles et al., 2021), particle MCMC has been employed to reconstruct the COVID-19 epidemic’s trajectory in Ireland. However, it is worth noting that while this method efficiently tracks observed data, it remains an offline approach and may not be suitable for real-time decision-making when rapid responses are required.
3.2.1 Model
We present a fully discrete stochastic extended-SEIR model to study the spread of COVID-19, with a structure tailored to focus on the change in transmission dynamics during the pandemic. In this model, individuals transition through compartments representing different stages of infection. The model comprises the following compartments: $S_{t}$ for susceptible individuals, $E_{t}$ for exposed but not yet infectious individuals, $A_{t}$ for asymptomatic infected individuals, $I_{t}$ for symptomatic infected individuals, and finally, $R_{t}$ is the count of individuals recovering or succumbing to the virus. The following discrete equations describe the dynamics of the model:
$S_{t+\delta t}$ $=S_{t}-Y_{SE}(t),$
$E_{t+\delta t}$ $=E_{t}+Y_{SE}(t)-Y_{E}(t),$
$A_{t+\delta t}$ $=A_{t}+Y_{EA}(t)-Y_{AR}(t),$
$I_{t+\delta t}$ $=I_{t}+Y_{E}(t)-Y_{EA}(t)-Y_{IR}(t),$ (17)
$R_{t+\delta t}$ $=R_{t}+Y_{AR}(t)+Y_{IR}(t),$
where $Y_{AB}(t)$ represents the number of individuals moving from compartment $A$ to compartment $B$ at time $t$ and defined by the following binomial distributions:
$Y_{SE}(t)\sim\mathrm{Bin}\left(S_{t},1-\exp\left(-\beta_{t}\frac{(I_{t}+r_{A}A_{t})}{N}\cdot\delta t\right)\right),$
$Y_{E}(t)\sim\mathrm{Bin}\left(E_{t},1-e^{-\sigma\cdot\delta t}\right),\quad Y_{EA}(t)\sim\mathrm{Bin}\left(Y_{E}(t),p_{A}\right),$ (18)
$Y_{AR}(t)\sim\mathrm{Bin}\left(A_{t},1-e^{-\gamma\cdot\delta t}\right),\quad Y_{IR}(t)\sim\mathrm{Bin}\left(I_{t},1-e^{-\gamma\cdot\delta t}\right).$
The model assumes a homogeneously mixing population of constant size:
$N=S_{t}+E_{t}+A_{t}+I_{t}+R_{t}.$ (19)
The models presented in (3.2.1)-(3.2.1), do not explicitly incorporate vaccination. This is because they are primarily designed to focus on the spread of the virus and the impact of various non-pharmaceutical interventions during a period when vaccines are either unavailable or not yet widely deployed. However, the effect of vaccination can easily be introduced by adding a compartment for vaccinated individuals and adjusting the transmission dynamics to reflect partial or complete immunity. Allowing the transmission rate to be time-varying seems reasonable since, throughout the outbreak, the intensity of disease transmission will tend to vary due to individuals’ behavioral changes or government lockdown measures. Hence, we consider a geometric random walk (RW) model for the transmission rate:
$\beta_{t}=e^{\log(\beta_{t-1})+w_{t}},\quad w_{t}\sim\mathcal{N}(0,\nu_{\beta}^{2}),$ (20)
where $\nu_{\beta}$ is the parameter controlling the innovation in the transmission rate, consistent with previous study of COVID-19 in Ireland (Cazelles et al., 2021). Giving the distribution of $\beta_{t}$, we can sequentially estimate the distribution of the effective reproduction number ($R_{\text{eff}}(t)$) that refers to the average number of secondary infections caused by a single infected individual during their infectious period in a population where some individuals are immune or other interventions have been implemented:
$R_{\text{eff}}(t)=\beta_{t}\left(\frac{(1-p_{A})+p_{A}\cdot r_{A}}{\gamma}\right).\frac{S_{t}}{N}.$ (21)
In general, when $R_{\text{eff}}(t)>1$, the number of observed cases will increase, while it will decrease if $R_{\text{eff}}(t)<1$. Consequently, policymakers can determine whether to relax or strengthen control measures based on whether $R_{\text{eff}}$ falls below the self-sustaining threshold of $1$.
3.2.2 Data and model fitting
The data source used to make inferences was restricted to daily confirmed cases counts extracted from the updated database of the Health Protection Surveillance Centre (HPSC) (https://COVID19ireland-geohive.hub.arcgis.com/). To integrate these observations into our model, we define two variables: $C^{I}_{t}$ representing the cumulative number of infectious cases. These are given by the following equation:
$C^{I}_{t+1}=C^{I}_{t}+Y_{E}(t)-Y_{EA}(t).$ (22)
Hence, $\mu_{t}=C^{I}_{t+1}-C^{I}_{t}$ represents the daily expected number of cases. Since daily reported case counts are subject to observation errors and typically exhibit overdispersion (where the conditional variance exceeds the conditional mean), we model the observed daily cases, $Z^{I}_{t}$, using a normal approximation (chosen for computational efficiency, Funk et al. (2018)) of the negative binomial distribution. The negative binomial distribution is a generalization of the Poisson distribution that accounts for overdispersion in the observation. Specifically, for the negative binomial distribution, the mean is $\mu_{t}$ and the variance is $\mu_{t}(1+\phi\mu_{t})$, where $\phi$ is the overdispersion parameter. To avoid singularities when $\mu_{t}=0$, we followed the approach of Funk et al. (2018), rounding variances smaller than 1 up to 1. Thus, the observed daily case counts follow a negative binomial distribution:
$Z^{I}_{t}\sim\mathcal{N}(\mu_{t},\mu_{t}(1+\phi\mu_{t})).$ (23)
We estimate the spread of COVID-19 in Ireland from February 29th, 2020, to February 28th, 2021, using the daily new infected count data of 365 days. The Irish population in 2020 was estimated to be $N=4,965,439$. The O-SMC2 uses $N_{\theta}=1000$ parameter particles, $N_{x}=500$ state particles, $M=5$ successive PMMH moves and a window size of $t_{k}=80$ days. Table C.1 gives the description and prior of the inferred parameters used in the model (3.2.1). These priors are based on previous studies of COVID-19 dynamics in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021). For the choice of prior distributions for the hyperparameters, $\nu_{\beta}$ (which governs the variability in the transmission rate) and $\phi$ (the overdispersion parameter), we adopt settings that allow for rapid changes in the transmission rate, ensuring the model can adapt to sudden shifts in disease dynamics, as discussed in (Funk et al., 2018). Regarding initial states, we set: $S_{0}=N-E_{0}-A_{0}$, $E_{0}=1$, $A_{0}\sim\mathcal{U}(\{10,\dots,50\})$ and $I_{0}=R_{0}=0$. The execution time for the 365 days considered here was less than 5 hours, making it suitable for practical real-time applications.
Figure 5: Model fitting with COVID-19 data in Ireland. The top row shows the estimated daily number of new cases. The observed data used for model fitting are shown as orange dots, while the red dots represent observed data that were excluded from fitting and used solely for forecast validation. The bottom row presents the weekly average of the time-dependent effective reproduction number. The solid black lines show the median estimate; shaded blue regions indicate posterior uncertainty quantiles. The vertical grey dashed lines represent the different government interventions. The vertical dashed violet line marks the last date of observations, beyond which the model is used to forecast the next 21 days.
Figure 5 illustrates the model’s fit to the daily number of COVID-19 cases reported in Ireland. The top plot shows the daily new infections, with three noticeable peaks: one in mid-April 2020, another in mid-November 2020, and a larger peak in early January 2021. The model successfully captures these three waves, with observed data (represented by the orange dots) falling within the 95% credible interval on most days. However, a sudden drop in reported cases around mid-January 2021 is not fully captured by the model. This anomaly could be due to the smoothing effect of Gaussian noise in the model, which limits its ability to promptly detect abrupt changes. Additionally, the drop may also reflect real-world factors such as reporting delays, testing changes, or a shift in public health measures that were not directly incorporated into the model. Beyond March 31, 2021, 2021, the model projects cases forward for 21 days, with data from April 1, 2021 (depicted as red dots) excluded from the fitting process. As shown in Figure 5, the observed data from April 1, 2021, fall within the 95% credible interval of the forecasted range, indicating that the model’s predictions align well with the actual trend.
The bottom plot in Figure 5 shows the weekly-averaged effective reproduction number, $R_{\text{eff}}(t)$, with corresponding daily estimates (not averaged) presented in Figure C.1. As anticipated, $R_{\text{eff}}(t)$ fluctuates in response to changes in the epidemic dynamics. The impact of the first lockdown is clearly evident, as it drives $R_{\text{eff}}(t)$ below one, effectively ending the first wave. A gradual increase in $R_{\text{eff}}(t)$ occurred in June 2020, due to the easing of restrictions. $R_{\text{eff}}(t)$ remained above one until mid-October, corresponding to the second wave of the epidemic. The second round of lockdown measures caused $R_{\text{eff}}(t)$ to decline by November 2020, falling below one once more. The emergence of the more transmissible Alpha variant, combined with increased social interactions during the holiday season, resulted in a sharp rise in cases toward the end of 2020. This was followed by a third round of lockdown measures, which once again brought $R_{\text{eff}}(t)$ below one (see the supplementary material of Jaouimaa et al. (2021) for a description of the individual lockdown).
In Figure 6, we present the sequential estimation of the parameters of model (3.2.1) over 365 days of the pandemic. The graphs display the posterior medians of the parameters, accompanied by their respective uncertainty intervals at each time step. The posterior distributions of the parameters obtained at the final time step (see Table C.1) are aligned with the results found in (Cazelles et al., 2021) using PMCMC methods.
Figure D.1 illustrates the application of the O-SMC2 for real-time monitoring, using partial datasets spanning four distinct time periods: February 29, 2020, to May 30, 2020; May 30, 2020, to August 29, 2020; August 29, 2020, to November 28, 2020; and November 28, 2020, to February 27, 2021. At each stage, the algorithm utilizes the results from the previous O-SMC2 run as prior information to compute the joint posterior $p(x_{i:j},\theta\mid y_{i:j})$. This sequential approach enables the efficient integration of new data while maintaining computational efficiency. Moreover, it significantly improves the methodology described in Dureau et al. (2013), which relied on PMCMC and required the model to be rerun from scratch for each new batch of data.
Figure 6: Filtered estimate of the parameter in the COVID-19 model using O-SMC2 algorithm. The solid black lines represent the median estimate; shaded blue regions indicate posterior uncertainty quantiles.
4 Conclusion
In this paper, we explored the application of an online variant of the Sequential Monte Carlo Squared (O-SMC2) method for real-time inference in stochastic epidemic models. This approach differs from standard SMC2 by integrating a mechanism that approximates the posterior distribution using a fixed window of recent observations, thereby reducing computational demands while maintaining parameter estimation accuracy. This adaptation makes O-SMC2 particularly suitable for dynamic disease modeling, where timely responses are critical. We demonstrated the effectiveness of O-SMC2 in the stochastic SEIR compartment model, validating its performance on both synthetic simulations and real-world COVID-19 data from Ireland. Our results show that O-SMC2, even with a modest window size, effectively tracks epidemic progression and generates robust online estimates of both static and dynamic epidemiological parameters. Moreover, the sequential nature of this method enables the efficient updating of estimates by using results from previous runs as prior information, eliminating the need to rerun the model from scratch as new data batches become available. This feature underscores the practicality of O-SMC2 for real-time epidemiological surveillance, enabling public health authorities to adapt interventions based on the latest information. However, we also acknowledge certain challenges, particularly in the early stages of disease transmission when available data may be sparse and less informative. This can hinder accurate parameter estimation during the onset of an outbreak. To address these limitations, future work will focus on refining our approach by integrating diverse modeling techniques that can better handle early-stage uncertainties and improve prediction accuracy.
Acknowledgements
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 21/FFP-P/10123.
Appendix A Approximation of the acceptance probability
During the rejuvenation step in the algorithm 2, when a new parameter $\theta^{*}$ is proposed at time $t$, an unbiased estimate of the marginal likelihood is computed by running Algorithm 1 over observations from time 1 to $t$. However, as time progresses, the computational cost of this procedure grows, making it less feasible for long-time series data. This limitation can be alleviated by considering only the most recent observations (Vieira, 2018).
Let the window size be denoted by $t_{k}>0$. By the law of total probability, the marginal likelihood at time $t$ can be factorized as:
$p(y_{1:t}|\theta)$ $=\left(\prod_{s=1}^{t-t_{k}-1}p(y_{s}|y_{1:s-1},\theta)\right)\times\left(\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta)\right)$
$=p(y_{1:t-t_{k}-1}|\theta)\times\left(\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta)\right).$ (24)
For $s\geq t-t_{k}$, if we assume that for a sufficiently large $t_{k}$, the older observations $y_{1:t-t_{k}-1}$ have minimal impact on the filtering estimate of the latent state at time $s-1$, i.e., $p(x_{s-1}|y_{1:s-1},\theta)\approx p(x_{s-1}|y_{t-t_{k}:s-1},\theta)$, then, from the result in (Kantas et al., 2015), we have:
[…]
$=p(y_{s}|y_{t-t_{k}:s-1},\theta).$ (25)
This approximation yields to:
$p(y_{1:t}|\theta)\approx p(y_{1:t-t_{k}-1}|\theta)\times p(y_{t-t_{k}:t}|\theta),$ (26)
where
$p(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{t-t_{k}:s-1},\theta)\approx\prod_{s=t-t_{k}}^{t}p(y_{s}|y_{1:s-1},\theta).$ (27)
Given the current $\theta$-particle, the new parameter $\theta^{*}$ at time $t$ is drawn from the region where most of the probability mass is located (Chopin, 2002). Under the assumption that the target parameter vector remains constant over time, and provided that the window size $t_{k}$ is large enough to capture the necessary temporal variability in the observations $y_{t-t_{k}:t}$, it is reasonable to assume that the historical likelihoods $p(y_{1:t-t_{k}-1}|\theta)$ and $p(y_{1:t-t_{k}-1}|\theta^{*})$ are approximately equal. Consequently, their contribution to the likelihood ratio becomes negligible. Under this assumption, the full likelihood ratio simplifies to:
$\dfrac{p(y_{1:t}|\theta^{*})}{p(y_{1:t}|\theta)}\approx\dfrac{p(y_{t-t_{k}:t}|\theta^{*})}{p(y_{t-t_{k}:t}|\theta)}.$ (28)
Thus, using the SMC estimate of the incremental likelihood for a fixed number of particles $N_{x}$ (see Equation (8)), we estimate the windowed likelihood $p(y_{t-t_{k}:t}|\theta)$ as:
$\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)=\prod_{s=t-t_{k}}^{t}\left(\frac{1}{N_{x}}\sum_{i=1}^{N_{x}}w_{s}^{i}(x_{s-1}^{i},x_{s}^{i})\right),$ (29)
and the acceptance probability for the PMMH kernel is given by:
$\alpha=\min\left\{1,\dfrac{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta^{*})p(\theta^{*})}{\hat{p}_{N_{x}}(y_{t-t_{k}:t}|\theta)p(\theta)}\times\dfrac{q(\theta|\theta^{*})}{q(\theta^{*}|\theta)}\right\}.$ (30)
Appendix B Posterior estimate in the simulation study
Here, we present the posterior estimates of model parameters from two experimental setups using the O-SMC2 algorithm. The experiments test varying window sizes $t_{k}$ to evaluate the computational efficiency and the impact of window length on parameter inference. For each window size, we report the posterior median and 95% credible intervals, comparing them to SMC2 estimates to assess both inference accuracy and computational savings. In our analysis, we define “Relative Efficiency ” as the ratio of the SMC2 computational time to the O-SMC2 computational time, providing a measure of the time reduction achieved by O-SMC2. This ratio quantifies the improvement in computational efficiency for each $t_{k}$ compared to SMC2, with larger ratios indicating greater computational efficiency.
Table B.1: Summary of parameters in Experiment 1. Posterior median and 95% credible intervals using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency ” denotes the computational time reduction factor compared to SMC2.
$t_{k}$ $\beta$ $\sigma$ $\gamma$ Relative Efficiency
20 0.611 (0.4751, 0.828) 0.389 (0.225, 0.652) 0.209 (0.135, 0.308) 1.70
40 0.603 (0.442, 0.870) 0.367 (0.219, 0.562) 0.213 (0.127, 0.338) 1.18
60 0.639 (0.499, 0.847) 0.334 (0.235, 0.447) 0.228 (0.158, 0.329) 1.10
80 0.638 (0.496, 0.845) 0.336 (0.235, 0.479) 0.224 (0.158, 0.325) 1.02
SMC2 0.594 (0.439, 0.841) 0.358 (0.235, 0.545) 0.210 (0.134, 0.343) -
Truth 0.6 0.333 0.2 -
Table B.2: Summary of parameters in Experiment 2. Posterior median and 95% credible intervals using O-SMC2 with varying window sizes $t_{k}$ at the final time step $T=100$. “Relative Efficiency ” denotes the computational time reduction factor compared to SMC2.
$t_{k}$ $\sigma$ $\gamma$ $\nu_{\beta}$ Relative Efficiency
20 0.477 (0.391, 0.574) 0.141 (0.130, 0.156) 0.136 (0.122, 0.148) 3.22
40 0.479 (0.391, 0.573) 0.141 (0.128, 0.154) 0.138 (0.125, 0.149) 1.59
60 0.476 (0.393, 0.571) 0.141 (0.132, 0.150) 0.139 (0.126, 0.149) 1.29
80 0.475 (0.387, 0.576) 0.142 (0.131, 0.154) 0.140 (0.127, 0.149) 1.06
SMC2 0.491 (0.406, 0.588) 0.143 (0.134, 0.153) 0.139 (0.127, 0.149) -
Truth 0.5 0.142 - -
While smaller $t_{k}$ values yield significant computational savings, this efficiency advantage decreases as $t_{k}$ increases. For most tested window sizes, the credible intervals align closely with those from SMC2, indicating minimal bias in this application. However, smaller windows may introduce minor biases if the window does not capture sufficient temporal dependencies, potentially resulting in slightly larger 95% credible intervals.
Figure B.1: Parameters’ posterior densities in Experiment 1 corresponding to different window sizes $t_{k}$.
Figure B.2: Parameters’ posterior densities in Experiment 2 corresponding to different window sizes $t_{k}$
It is important to note that the windowing mechanism is applied only during the resampling step (i.e., when ESS drops below 50% of the total number of parameter particles in our case). Thus, the greater the frequency of resampling at time $t$ (especially when $t$ is significantly larger than $t_{k}$), the more computationally beneficial O-SMC2 becomes. In Figure B.3, we observe that the frequency of resampling at times well beyond $t_{k}=20$ gradually decreases in Experiment 1, while it remains more frequent in Experiment 2. This explains the greater computational gain observed in Experiment 2.
Figure B.3: Effective Sample Size. The left graph shows the the evolution of the ESS in Experiment 1 and the right one the evolution in Experiment 2.
Appendix C Parameter settings in the COVID-19 model
Table C.1 summarizes parameters from the COVID-19 model (Equation (3.2.1)), including hyperparameters (Equations (20) and (23)). For each parameter, prior and posterior estimates (median and 95% credible interval) are provided. Priors were informed by COVID-19 dynamics literature in Ireland (Irish Department of Health, 2020; Cazelles et al., 2021).
[…]
Parameter Description Prior value Posterior median (95% CrI)
$r_{A}$ Reduction factor of transmission from $A_{t}$ $\mathcal{U}(0.1,0.5)$ 0.370 (0.304, 0.445)
$p_{A}$ Fraction of infected asymptomatic cases $\mathcal{U}(0.3,1)$ 0.413 (0.313, 0.491)
$\beta_{0}$ Initial condition transmission rate $\mathcal{U}(0.6,0.8)$
$\sigma$ Rate from exposed to infectious $\mathcal{TN}_{[1/3,1/5]}(1/4,0.1)$ 0.286 (0.256, 0.307)
$\gamma$ Removal rate $\mathcal{TN}_{[1/7.5,1/4.5]}(1/6,0.2)$ 0.184 (0.165, 0.213)
$\nu_{\beta}$ Volatility of the Brownian process $\mathcal{TN}_{[0.05,15]}(0.1,0.05)$ 0.137 (0.118, 0.149)
$\phi$ Overdispersion $\mathcal{U}(0.01,0.2)$ 0.100 (0.088, 0.114)
Figure C.1: Daily estimates of the time-dependent transmission rate and effective reproduction number. The solid black lines represent the median estimate of $R_{\text{eff}}(t)$, with shaded blue regions indicating posterior uncertainty quantiles. Weekly averaged estimates of $R_{\text{eff}}(t)$ are shown in Figure 5.
Appendix D Application in real-time surveillance
The methodology developed in this paper is applied to a real-time surveillance setting, demonstrating how our model can track evolving dynamics in sequential data. By integrating new observations over time, the model provides insights into the system’s behavior. The focus is on an adaptive framework that updates the posterior distribution, conditioned on incoming data, while utilizing results from previous O-SMC2 runs as prior information. The results are shown in Figure D.1, where we observe that there is no significant difference compared to the results obtained from running the model from scratch (see Figure 5).
Figure D.1: Real-time application of O-SMC2 to sequential data: (Row 1) February 29, 2020, to May 30, 2020; (Row 2) May 30, 2020, to August 29, 2020; (Row 3) August 29, 2020, to November 28, 2020; and (Row 4) November 28, 2020, to February 27, 2021. Each row depicts the sequential estimates of daily new cases and the weekly average effective reproduction number ($R_{\text{eff}}(t)$). The solid black lines represent the median estimate, while the shaded blue regions indicate posterior uncertainty quantiles.